A voltage V across a resistance R generates a current I=V/R. If a constant voltage of 8 volts is put across a resistance that is increasing at a rate of 0.3 ohms per second when the resistance is 3 ohms, at what rate is the current changing? (Give units.)
1
Expert's answer
2020-03-02T09:33:49-0500

Given

V = 8 volts ( constant )

Resistance is increasing at the rate of 0.3 ohms per second

i.e. "\frac{dR}{dt}=0.3\:" ohms/sec

We need to find rate of change of current when resistance is 3 ohms

"\frac{dI}{dt}=?" when R= 3 ohms


We know the relation

"I=\frac{V}{R}"

Differentiating with respect to time

"\frac{d}{dt}\left(I\right)=\frac{d}{dt}\left(\frac{V}{R}\right)"

since voltage is constant, it can be taken out.

"\frac{dI}{dt}=V\cdot \frac{d}{dt}\left(\frac{1}{R}\right)"

"\frac{dI}{dt}=V\cdot \:\left(-\frac{1}{R^2}\cdot \frac{dR}{dt}\right)"

substituting values

"\frac{dI}{dt}=8\cdot \:\left(-\frac{1}{3^2}\cdot 0.3\right)=-\frac{4}{15}" amperes per second

Negative sign indicates current is decreasing with respect to time.

So current is decreasing at the rate of 4/15 amperes per second when R is 3 ohms.